A firm is operating in the short run. Here is some of the information about the firm's operation.

  1. The short run fixed costs for the firm are $50.
  2. The wage rate for each employee is $120 per day
  3. The Production Function is below

Production Function

Labor

Daily Output

1

60

2

130

3

200

4

260

5

310

6

320

7

325

8

326

Variable Resources

Output

MP

TFC

TVC

TC

MC

ATC

AFC

AVC

0

0

50

0

50

1

60

60

50

120

170

2

2.83

.83

2

2

130

70

50

240

290

1.71

2.23

.38

1.85

3

200

70

50

360

410

1.71

2.05

.25

1.8

4

260

60

50

480

530

2

2.04

.19

1.85

5

310

50

50

600

650

2.4

2.10

.16

1.94

6

320

10

50

720

770

12

2.41

.16

2.25

7

325

5

50

840

890

24

2.73

.15

2.58

8

326

1

50

960

1010

120

3.1

.15

2.94

Using the chart , answer the question.

-Under economic reasoning, should the firm produce if the price temporarily drops to $1.83? Why or why not?
Answer #1

A competitive firm can increase its output level as long as MC is lower than or equal to Price.

We can see that MC<P for variable resource=3 (output 200 units)

So, optimal output should be 200 units.

Now let us see other factors

ATC at this level=$2.05

Price is less than ATC. It means that firm will make a loss at this output.

AVC at this level=$1.80

Price is higher than AVC at this level. It means that firm is able to recover its variable expenses as well as some of the fixed cost.

Profit=TR-TC=1.83*200-410=-$44

Firm will make a loss of $44 if it continues production and produces optimal output.

Now if firm shuts down in short run,

Profit=TR-TC=0-50=-$50

Firm will make a loss of $50 if it shuts down operations in short run.

We can see that loss is higher if firm shuts down in short run. So, firm should continue its operations in short run to minimize losses.